OpenStudy (kaylala):

trigonometry: Identify if the underlined word/s is true or false If false change it to the appropriate one A central angle of 114.6deg in a circle of radius 1 unit intercepts an arc of equal to 114.6units 114.6units is the underlined word

OpenStudy (***[isuru]***):

convert 114.6 deg in to radiants

OpenStudy (kaylala):

is it 19.1π/30?

OpenStudy (***[isuru]***):

yes!

OpenStudy (***[isuru]***):

now use \[S = r \theta\] :P

OpenStudy (loser66):

no need to do that , it's true since radius =1 , it is false if radius \(\neq 1\)

OpenStudy (kaylala):

S=19.1π/30 ? is it?

OpenStudy (***[isuru]***):

yes ... i agree loser66 ... but it;s the formal method... so i thought it would be better to say it first

OpenStudy (kaylala):

i dont know what is its unit of this measure?

OpenStudy (***[isuru]***):

@kaylala it is exactly as loser66 said.. since the raius is 1 unit... the arc is equal to the angle..

OpenStudy (kaylala):

so my answer is in degree? 114.6deg?

OpenStudy (***[isuru]***):

u dont need to know the unit of the radius... because the unit cancel out since the radius and the arc length both belong to the same unit

OpenStudy (kaylala):

i see

OpenStudy (kaylala):

but s=114.6units. not 114.6deg are those equal still?

OpenStudy (kaylala):

is 114.6units equal to 114.6deg?

OpenStudy (***[isuru]***):

No. but the question says the angle is 114.6 degrees and the arc is 114.6units

OpenStudy (***[isuru]***):

it doesnt say 114.6 angle is equal to 114.6 unit

OpenStudy (kaylala):

wait i'm confused

OpenStudy (loser66):

the formula is the length of the arc L = the angle * the radius. your radius = 1 if its unit is cm, then, r = 1 cm, right? if its unit is m, then r = 1m, right? If they don't use the unit of measurement like m, cm, km...., but "unit" then, the radius is 1 unit, right? so??? the length of the arc L = 114.6 * 1 unit = 114.6 unit

OpenStudy (kaylala):

hence it's false? @Loser66 arc L is 114.6 unit and not 114.6deg???

OpenStudy (kaylala):

@ isuru?

OpenStudy (***[isuru]***):

no it;s not false ....

OpenStudy (***[isuru]***):

the angle has a unit ... so it's in degrees

OpenStudy (kaylala):

oh i get it now

OpenStudy (kaylala):

thanks :)

OpenStudy (***[isuru]***):

u r welcome!